# Pre-Processing Matters: Similarity-based Simulated Annealing Approach for Fast Parameter Tuning in WSI Classification

### v2

@@ Abstract
Pre-processing whole slide images (WSIs) can impact classification performance. Our study shows that using fixed hyper-parameters for pre-processing out-of-domain WSIs can significantly degrade performance. Therefore, it is critical to search domain-specific hyper-parameters during inference. However, searching for an optimal parameter set is time-consuming. To overcome this, we propose SSAPT, a novel Similarity-based Simulated Annealing approach for fast parameter tuning to enhance inference performance on out-of-domain data. The proposed SSAPT achieves 5% to 50% improvement in accuracy with ×$\times 5$ times faster parameter searching speed on average.
Introduction
@@
Following the success of early CPATH applications, dataset sizes have increased, prompting more multicentric efforts to address variability in staining, image quality, scanning characteristics, and tissue preparation across different labs. This has highlighted a known issue where CPATH algorithms perform best on data they were trained on but less well on data from other sources. Generalization continues to pose a major challenge in this field, as significant differences in clinical variables between tissue source sites can adversely affect the performance of histopathological tasks (Van der Laak, Litjens, and Ciompi 2021; Dehkharghanian et al. 2023; Howard et al. 2021).
@@
State-of-the-art MIL models have shown promising improvements in WSI classification on out-of-domain data by training on large, diverse datasets (Zhang et al. 2022; Campanella et al. 2019; Lu et al. 2021; Shao et al. 2021; Li, Li, and Eliceiri 2021; Tang et al. 2023). Out-of-domain classification occurs when a model is trained on one dataset and tested on another from a different domain. For instance, a model might be trained using Camelyon16 data (Litjens et al. 2018) and then evaluated using data from various centers in Camelyon17. However, these studies typically only report average accuracy, neglecting the variability in performance across different centers. As indicated in Table 1, the performance of a model trained on Camelyon16 can vary significantly at specific centers within Camelyon17.
Figure 1: Different pre-processing methods for Whole Slide Images (WSIs) generate distinct features, which in turn affect inference performance. Retaining excessive background and discarding too much tissue can result in poor inference outcomes.
We discovered that unsatisfactory performance is often due to using the same inference hyper-parameter across different centers. In real-world scenarios, we apply the fixed pre-processing parameters to extract features and evaluate performance on new sub-datasets. However, this approach typically results in poor out-of-domain performance (Howard et al. 2021). For example, MIL models (Lu et al. 2021; Cui et al. 2023; Li, Li, and Eliceiri 2021) have extremely low accuracy at specific centers within TCGA and Camelyon17 datasets. Conversely, tailoring optimal pre-processing parameters for each center significantly improves out-of-domain performance (as illustrated in Figure 2).
Determining optimal pre-processing parameters for histopathological tasks is challenging due to the vast size of Whole Slide Images (WSIs) and the significant computational resources needed. Traditional grid search methods for datasets like Camelyon or TCGA are inefficient, taking several hours to evaluate one parameter group. The whole hyper-parameter space contains hundreds of thousands of possible combinations. Therefore, quicker hyper-parameter search techniques are essential for histopathological tasks.
@@
Figure 2: The performance of out-of-domain inference varies with preprocessing parameters across various MIL models and datasets. Consequently, we suggest that each specific center within the dataset should adopt its own preprocessing parameters to maximize performance. The original method involves all centers using fixed default preprocessing hyperparameters, whereas the optimal method allows each center to use its own specific preprocessing hyperparameters determined by our proposed SSAPT.
We have observed that varying pre-processing methods can significantly impact model performance during both inference and training. This paper focuses on how pre-processing parameters influence inference performance, particularly with out-of-domain data. This paper proposes SSAPT, Similarity-based Simulated Annealing for Hyper-parameter Tuning, a method that improved the accuracy of inferring out-of-domain data by 5% to 50% across various MIL models and datasets. Previous studies, such as those cited in (Choi et al. 2024), only explored hyper-parameter optimization in terms of per-layer learning rates and loss weight coefficients , without tailoring their approaches to histopathology applications. The key contributions of this paper are:
@@
  • We present SSAPT, Similarity-based Simulated Annealing for fast and effective parameter tuning. This algorithm enhances inference performance at Camelyon 17’s center 1, boosting accuracy from 0.512 to 0.847;
  • We expand the proposed SSAPT to include other MIL models using various public datasets, achieving improvements in accuracy ranging from 5% to 50% for out-of-domain data across multiple MIL models and datasets;
  • The proposed SSAPT is the first fast hyper-parameter search specifically designed for histopathological tasks.
Related work
MIL models. DSMIL and CLAM employ instance-loss for training, with DSMIL featuring an instance-level classifier. Bayes-MIL focuses on patch-level uncertainty. TransMIL (Shao et al. 2021) treats patches as tokens to utilize transformer architecture. MHIM (Tang et al. 2023) employs a masked hard instance mining strategy. All these models are designed at the instance/patch level, which leads to poor out-of-domain performance. In contrast, ABMIL (Ilse, Tomczak, and Welling 2018) operates solely at the bag level, resulting in lower accuracy in in-domain performance across various datasets but achieving greater robustness. However, it is impossible to completely avoid instance-level design in developing new MIL models for histopathology.
Pre-processing of WSI in histopathology. Variations in tissue processing techniques, including chemical fixation or freezing, dehydration, embedding, and staining, can alter the visual characteristics of the tissue slide in ways that are both non-uniform and non-linear. These changes occur across different tissue types and laboratories (Asif et al. 2023; Yagi 2011), consequently impact the performance in deep learning. Salvi et al.; Gurcan et al. summize the pre-processing impact for WSI analysis on deep learning frameworks. Tellez et al. quantifies the effects of data augmentation and stain color normalization. (Lee et al. 2022)
There are several research works that specifically explore the issue of dataset bias in CPath. Pocevičiūtė et al. quantifies the impact of domain shift in attention-based MIL and point out that MIL performance is worse than what is reported using in-domain test data. Dehkharghanian et al.; Howard et al. discusses bias on histopathological data. Important clinical variables have been shown to be significantly different between tissue source sites, affecting the generalization issue for histopathological tasks.
Hyper-parameters search and optimization in machine learning. Falkner, Klein, and Hutter propose a Bayes-optimization with Hyperband strategy and evaluate it in small datasets. Choi et al. search hyper-parameters in per-layer learning rates and loss weight coefficients. These previous works are not designed for histopathology.
Motivation and Observation
@@
Public datasets often combine data from various domains (Torralba and Efros 2011; Pocevičiūtė et al. 2023), obscuring true performance metrics within specific areas. Studies (Dehkharghanian et al. 2023; Howard et al. 2021; Hosseini et al. 2024) show that domain-specific variations in digital histology can affect the accuracy and bias of deep learning models. In practice, these variations necessitate frequent retraining of models due to changing data domains, which is impractical due to high computational costs. Thus, improving model performance without additional training is essential.
Pre-processing Impacts Inference Performance for Different Domains.
@@
We found there exists significant variability in inference performance on out-of-domain data across various methods (refer to Table 1). For instance, the CLAM model trained on the Camelyon16 dataset achieves an overall inference accuracy of 0.86 on the Camelyon17 dataset but drops to 0.512 at Center 1 of Camelyon17, indicating inconsistent performance. We have also tested our approach in other centers and datasets, as detailed in the Appendix.
We employ a grid search to fully understand how pre-processing parameters influence outcomes. We discovered that modifying these parameters affects out-of-domain (OOD) inference accuracy differently across centers. The default pre-processing setting yields varying results depending on the center as shown in Fig 3, where identical hyper-parameters may perform well in one center but poorly in another. Similar variations are observed across other datasets and MIL models (additional details provided in Appendix). This inconsistency stems from instance/patch-level design choices. We conduct experiments to examine how adjusting bag loss and instance loss weights for DSMIL and CLAM impacts OOD inference performance. Comprehensive details about these instance-level design experiments are available in the Appendix.
Inefficient Manual Pre-Processing Parameter Search.
It’s beneficial to optimize parameter performance, but finding the best settings is time-consuming; it takes approximately 7 hours to extract features and evaluate a set of parameters on an RTX3080 (details in Appendix). Therefore, we need more efficient parameter search methods. Common strategies like grid search require excessive computational resources and time, rendering them impractical for this scenario.
@@
  • Can modifying only the preprocessing improve inference performance on out-of-domain data?
@@
  • Do these observations apply to other state-of-the-art (SOTA) models?
@@
Therefore, we introduce our approach in fast hyper-paramater search in pre-processing parameters of histopathology.
@@
Fast Hyper-Parameter Search for Inference Performance in Out-Of-Domain Data
@@
In this section, we formalize hyperparameter optimization and present our SSAPT algorithm step by step. The proposed SSAPT algorithm includes PSNR-based mechanism for limiting the current searching space in each iteration and Simulated Annealing to avoid local optima.
Definition of Problem
A good choice of hyperparameters can significantly improve out-of-domain inference performance, yet optimal hyperparameters vary, depending on the datasets and the models. Formally, we denote the fast hyper-parameter search algorithm as $H_{opt}=SSAPT(H,D_{val}^{n})$ and its hyper-parameters as $H\in\Phi$, where $\Phi$ is the hyper-parameter space. We denote the training and validation datasets as $D_{train}$ and $D_{val}$, respectively. The model $M$ is trained in $D_{train}$, where all WSIs are using fixed hyper-parameter $H_{train}$. We denote the next inferring processing as $y=Infer(H_{i},D_{val}^{n})$. Hyperparameter optimization finds hyperparameters that maximize some performance metric $y$. We formalize hyperparameter optimization as
$H_{n}^{*}=\underset{H\in\Phi}{\operatorname{arg\,max}}\,\text{Infer}(SSAPT(H,D_{val}^{n}),D_{val}^{n})$ (1)
General Assumptions
@@
We assume access to two types of datasets: (1) a large dataset $D_{train}$ for training to produce the foundation model subsequently used to test out-of-domain performance. (2) Multiple subdataset $\{D_{val}=D_{val}^{1},D_{val}^{2},\cdots,D_{val}^{n}\}$. The validation sub-dataset $D_{val}$ is much smaller than $D_{train}$ and is used only for optimization at the outer level, not for fine-tuning. Note that these test distributions are distinct and are never shown to the model before the final evaluation.
Let $M$ denote a pre-trained foundation model trained in a large dataset $D_{train}$ with the pre-processing parameter $H_{train}$. Note that the model $M$ is trained by k-fold, StratifiedKFold, or Monte Carlo cross-validation, so $D_{train}$ actually contains a sub-dataset of training, validation, and testing while training.
@@
Now, there are many out-of-domain datasets $\{D_{val}=D_{val}^{1},D_{val}^{2},\cdots,D_{val}^{n}\}$. They are pre-processed by hyper-parameter $H_{1},H_{2},\cdots,H_{n}$ to obtain the features $F_{val}^{1},F_{val}^{2},\cdots,F_{val}^{n}$ for inferring. The out-of-domain performance of model $M$ would be tested by the features $F_{val}^{n}$ that are preprocessed by the hyper-parameter $H_{n}$. We adapt our algorithms as $H_{n}^{*}=SSAPT(H_{n},D_{val}^{n}$) to search the optimal hyper-parameters $H_{n}$ for out-of-domain dataset $D_{val}^{n}$.
In our experiments, the performance metric $y$ is accuracy, due to the unbalanced datasets in histopathological tasks. For instance, some centers contain only one category of WSI in TCGA-NSCLC, resulting in the situation that AUC and F-score cannot be calculated. Furthermore, extremely low accuracy is reported in some challenging cases, leading to the situation that accuracy is lower than 0.5, but the AUC is still higher than 0.8 (as shown in Table 1).
Pre-Processing Problem Modeling
@@
We define the task of identifying the optimal hyper-parameter for enhanced out-of-domain performance as an optimization problem, as described in Equation 1. In practice, this involves repeatedly applying different features within the same foundational model using various hyperparameters $H_{n}^{i}$. This process helps us determine the most effective final hyper-parameters $H_{n}$ for each validation center $D_{val}^{n}$. We assess the effectiveness of each $H_{n}^{i}$ by evaluating its performance on the OOD validation set $D_{val}^{n}$.
@@
Next we consider the optimization problem for histopathological tasks. The optimization of $H_{n}$ is actually the optimization of a combination of pre-processing parameters, not a single variable. Previous pre-processing of WSIs encompasses a series of standard procedures. Initially, tissue segmentation is executed automatically on each slide at a reduced magnification. This involves generating a binary mask for the tissue regions by applying a binary threshold to the saturation channel of the downsampled image, following its conversion from RGB to HSV color space. Morphological closing and median blurring are employed for smoothing the contours of the detected tissue and in minimizing artifacts. Subsequently, the approximate contours of both the tissue and the tissue cavities are assessed and filtered based on their area, culminating in the creation of the final segmentation mask. We complete set of pre-processing hyperparameters for optimization in histopathological tasks is thus $H_{n}=[x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}]$, where each $x_{i}$ has its own hyper-parameter space.
Optimization of Fast Hyper-Parameter Search
Good Pre-Processed WSIs show High Structural Similarity.

### v3

# BAHOP: Similarity-based Basin Hopping for A fast hyper-parameter search in WSI classification

@@ Abstract
Pre-processing whole slide images (WSIs) can impact classification performance. Our study shows that using fixed hyper-parameters for pre-processing out-of-domain WSIs can significantly degrade performance. Therefore, it is critical to search domain-specific hyper-parameters during inference. However, searching for an optimal parameter set is time-consuming. To overcome this, we propose BAHOP, a novel Similarity-based Basin Hopping optimization for fast parameter tuning to enhance inference performance on out-of-domain data. The proposed BAHOP achieves 5% to 30% improvement in accuracy with ×$\times 5$ times faster on average.
1 Introduction
@@
Following the success of early computational pathology applications, dataset sizes have increased, prompting more multicentric efforts to address variability in staining, image quality, scanning characteristics, and tissue preparation in different laboratories. This has highlighted a known issue where computational pathology algorithms perform best on data on which they were trained but less well on data from other sources. Generalization continues to pose a major challenge in this field, as significant differences in clinical variables between tissue source sites can adversely affect the performance of histopathological tasks [26, 6, 11].
@@
State-of-the-art multiple instance learning (MIL) models have shown promising improvements in whole slide image (WSI) classification on out-of-domain (OOD) data by training on large, diverse datasets [31, 2, 17, 21, 15, 23]. Out-of-domain classification occurs when a model is trained on one dataset and tested on another from a different domain. For instance, a model might be trained using Camelyon16 data [16] and then evaluated using data from various Camelyon17 centers. As indicated in Table 1, the performance of a model trained in Camelyon16 can vary significantly at specific centers within Camelyon17.
@@
Figure 1: The performance of out-of-domain inference varies with preprocessing parameters across various MIL models and datasets. Consequently, we suggest that each specific center within the dataset should adopt its own preprocessing parameters to maximize performance. The original method involves all centers using fixed default preprocessing hyperparameters, whereas the optimal method allows each center to use its own specific preprocessing hyperparameters determined by our proposed BAHOP.
Figure 2: PSNR-based Basin Hopping hyper-parameter optimization for out-of-domain inference in WSIs. The hyper-parameter optimization for the pre-processing task is a non-convex optimization that contains many large local optima. Our BAHOP is developed for fast search, and the jumping range of each iteration is controlled by the PSNR threshold.
We discovered that one factor that leads to unsatisfactory performance is the same inference hyper-parameter across different centers. We evaluate the fixed pre-processing parameters’ performance on several sub-datasets. The experiment result shows that fixed pre-processing parameters typically result in poor out-of-domain performance [11]. For example, MIL models [17, 5, 15] have extremely low accuracy at specific centers within the TCGA [29] and Camelyon17 [16] datasets using fixed parameters. Conversely, tailoring optimal pre-processing parameters for each center significantly improves out-of-domain performance (as illustrated in Figure 1).
Determining optimal pre-processing parameters for histopathological tasks is challenging. The parameter search space is over $10^{5}$ due to the huge size of giga-pixel Whole Slide Images (WSIs) and the significant computational resources needed. Previous studies [3] only explore hyper-parameter optimization in learning rates and loss weight coefficients without tailoring their approaches to histopathology applications. When adopting hyper-parameter tuning methods, traditional grid search methods for datasets like Camelyon or TCGA are inefficient, taking several hours to evaluate one parameter group. The whole hyper-parameter space contains hundreds of thousands of possible combinations. Therefore, quicker hyper-parameter search techniques are essential for histopathological tasks.
This paper focuses on how pre-processing parameters influence inference performance, particularly with out-of-domain data, and proposes an efficient parameter search method specifically designed for WSI classification. We propose a Similarity-based Basin Hopping (BAHOP) for Hyper-parameter tuning and improving the accuracy of inferring out-of-domain data across various MIL models and datasets (as shown in Fig. 1). The key contributions of this paper are:
@@
  • We present BAHOP, Similarity-based Basin Hopping for fast and effective parameter tuning. This algorithm enhances inference performance at Camelyon 17’s center 1, boosting accuracy from 0.512 to 0.846;
  • We expand the proposed BAHOP to include other MIL models using various public datasets, achieving improvements in accuracy ranging from 5% to 30% for out-of-domain data across multiple MIL models and datasets;
  • The proposed BAHOP is the first fast hyper-parameter search designed explicitly for histopathological tasks.
@@
2 Related Work
Pre-processing of WSI in histopathology.
Variations in tissue processing techniques, including chemical fixation or freezing, dehydration, embedding, and staining, can alter the visual characteristics of the tissue slide in ways that are both non-uniform and non-linear. These changes occur across different tissue types and laboratories [1, 30], consequently, impacting the performance of deep learning. Salvi et al. [20], Gurcan et al. [9], Wang et al. [28] summarize the pre-processing impact for WSI analysis on deep learning frameworks. Tellez et al. quantifies the effects of data augmentation and stain color normalization [14].
There are several research works that specifically explore the issue of dataset bias in computational pathology. Pocevičiūtė et al. [18] quantifies the impact of domain shift in attention-based MIL and points out that MIL performance is worse than what is reported using in-domain test data. Several works [6, 11] discuss bias in histopathological data. Important clinical variables have been shown to be significantly different between tissue source sites, affecting the generalization issue for histopathological tasks.
Hyper-parameters search and optimization in machine learning.
[8] propose a Bayes-optimization with Hyperband strategy and evaluate it in small datasets. [3] search hyper-parameters in per-layer learning rates and loss weight coefficients. These previous works are not designed for histopathology.
Hyper-parameter optimization in histopathological tasks is different than optimization problems in a linear function. The optimization in linear function or small machine learning model with a small dataset can be optimized in many iterations in a short time. However, the optimization in the machine learning model for histopathological tasks is totally different. WSIs are giga-pixel images that demand significantly higher computational resources. Despite advancements in hardware, memory constraints still pose a significant challenge in the digital processing of WSIs in current research [4, 7]. Additionally, nature images have very different data distribution compared to the whole slide image [19]. Therefore, designing a hyper-parameter optimization strategy for computational pathology is necessary. The proposed BAHOP algorithm is the first hyper-parameter optimization strategy designed for histopathological tasks.
3 Motivation and Observation
@@
Public datasets often combine data from various domains [25, 18], obscuring true performance metrics within specific areas. Studies [6, 11, 10] show that domain-specific variations in digital histology can affect the accuracy and bias of deep learning models. In practice, these variations necessitate frequent retraining of models due to changing data domains, which is impractical due to high computational costs. Thus, improving model performance without additional training is essential.
3.1 Pre-processing Impacts Inference Performance for Different Domains.
Figure 4: All the heatmaps circled by red boxes correspond to the default hyper-parameter and predict wrong, while heatmaps circled by blue boxes correspond to optimal hyperparameters with correct predictions. Fig.A: Hyper-parameter optimization starts from the default parameter, which is the same as the pre-trained model. The default pre-processed WSI drops many patches that get a high attention score in optimal pre-processed WSI. Fig.B, the dataset is TCGA-NSCLC where dropping tissue can improve accuracy. Fig.C: Some hyper-parameters drop the region of interest(ROI) during pre-processing.
@@
We found there exists significant variability in inference performance on out-of-domain data across various methods (refer to Table 1). We employ a grid search to fully understand how pre-processing parameters influence outcomes in center 1 of Camelyon 17 in this motivation experiment. For instance, the CLAM [17] model trained on the Camelyon16 dataset achieves an overall inference accuracy of 0.86 on the entire Camelyon17 dataset but drops to 0.512 at Center 1 of Camelyon17, indicating inconsistent performance. We also evaluate our approach in other centers across different datasets, as detailed in Fig 3 and Appendix. Similar variations are observed across other datasets and MIL models, and the out-of-domain inference is extremely poor in some cases. What’s more, the default pre-processing setting yields varying results depending on the center as shown in Fig 3, where identical hyper-parameters may perform well in one center but poorly in another.
Observation 1
The effection of the pre-processing parameters varies across different datasets. As shown in Fig. 4 and the Appendix, the detection of foreground contours, controlled by several hyperparameters, often results in the exclusion of large areas, which are treated as holes. For instance, adipocyte cells are frequently excluded. However, some tumor cells surrounded by fibers, adipocyte cells, or lipid droplets may also be filtered out due to preprocessing parameters and discarded. In Fig. 4(a) and (c), excessive patch removal leads to incorrect predictions, while in Fig. 4(b), it improves prediction accuracy when applied to a different dataset. This variability makes it difficult to quantify the impact of specific patches, especially since the number of patches for a single WSI in a MIL model can exceed $10^{5}$. While some studies, such as Javed et al. [13], address this issue, there is no consensus on how specific patches influence machine learning models in computational pathology. An adaptive algorithm is needed to address this challenge in histopathological tasks.
Observation 2
Pre-processing optimization is inherently a non-convex problem with numerous large local optima (as shown in Fig. 2). Our objective is to achieve an optimal result with reduced time and computational costs, where this optimum does not need to be a global one.
Rather than identifying specific cell types within tissue areas, we approach this histopathological task from a machine learning perspective. The pre-processing of WSIs focuses on segmenting tissue regions and discarding patches identified as holes or backgrounds. These patches are selected and filtered based on pre-defined parameters, while most tissue areas containing cells are retained. Consequently, the variations in inference performance caused by multiple interacting hyperparameters can be framed as a hyperparameter optimization problem. This problem can be effectively addressed using an adaptive algorithm tailored for this histopathological task.
3.2 Inefficient Manual Pre-Processing Parameter Search.
In real-world scenarios, retraining machine learning models for every new dataset is often impractical. While optimizing parameter performance is advantageous, finding the optimal settings is highly time-consuming. For instance, evaluating a single set of parameters, including feature extraction, takes approximately 7 hours on an RTX3080 (details in Appendix). Traditional optimization techniques, such as tuning learning rates, batch sizes, or loss weight coefficients, cannot be directly applied to histopathological tasks. Consequently, more efficient parameter search methods are required. Common strategies like grid search demand excessive computational resources and time, rendering them unsuitable for histopathological applications.
@@
  • Can we improve inference performance on out-of-domain data just by modifying the preprocessing parameters?
@@
  • Do these observations apply to other multiple instance learning (MIL) models?
@@
Therefore, we introduce our approach in a fast hyper-parameter search in pre-processing parameters of histopathology.
@@
4 Fast Hyper-Parameter Search for Inference Performance in Out-Of-Domain Data
@@
In this section, we formalize hyperparameter optimization and present our BAHOP algorithm step by step. The proposed BAHOP algorithm includes PSNR-based mechanism for limiting the current searching space in each iteration and Basin Hopping to avoid local optima.
4.1 Definition of Problem
@@
A good choice of hyperparameters can significantly improve out-of-domain inference performance, yet optimal hyperparameters vary, depending on the datasets and the models. We assume access to two types of datasets: (1) a large dataset $D_{train}$ for training to produce the foundation model subsequently used to test out-of-domain performance. (2) Multiple subdataset $\{D_{val}=D_{val}^{1},D_{val}^{2},\cdots,D_{val}^{n}\}$. The validation sub-dataset $D_{val}$ is much smaller than $D_{train}$ and is used only for optimization at the outer level, not for fine-tuning. Note that these test distributions are distinct and are never shown to the model before the final evaluation.
@@
Let $M$ denote a pre-trained foundation model trained in a large dataset $D_{train}$ with the pre-processing parameter $H_{train}$. Now, there are many out-of-domain datasets $\{D_{val}=D_{val}^{1},D_{val}^{2},\cdots,D_{val}^{n}\}$. They are pre-processed by hyper-parameter $H_{1},H_{2},\cdots,H_{n}$ to obtain the features $F_{val}^{1},F_{val}^{2},\cdots,F_{val}^{n}$ for inferring. The out-of-domain performance of model $M$ would be tested by the features $F_{val}^{n}$ that are preprocessed by the hyper-parameter $H_{n}$ as $y=Infer(H_{n},D_{val}^{n})$, where $y$ is the performance metric.
@@
We define the histopathological task of identifying the optimal pre-processing for enhanced out-of-domain performance as a hyper-parameter optimization problem, as described in Equation 1, where $\Phi$ is the hyper-parameter space. This process helps us determine the most effective final hyper-parameters $H_{n}^{*}$ for each validation center $D_{val}^{n}$.
$H_{n}^{*}=\underset{H\in\Phi}{\operatorname{arg\,max}}\,\text{Infer}(BAHOP(H,D_{val}^{n}),D_{val}^{n})$ (1)
4.2 Pre-Processing Problem Modeling
@@
Next, we consider the optimization problem for histopathological tasks. The optimization of $H_{n}$ is actually the optimization of a combination of pre-processing parameters, not a single variable. Previous pre-processing of WSIs encompasses a series of standard procedures. Initially, tissue segmentation is executed automatically on each slide at a reduced magnification. This involves generating a binary mask for the tissue regions by applying a binary threshold to the saturation channel of the downsampled image, following its conversion from RGB to HSV color space. Morphological closing and median blurring are employed for smoothing the contours of the detected tissue and minimizing artifacts. Subsequently, the approximate contours of both the tissue and the tissue cavities are assessed and filtered based on their area, culminating in the creation of the final segmentation mask.
We complete set of pre-processing hyperparameters for optimization in histopathological tasks is thus $H_{n}=[x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}]$, where each $x_{i}$ has its own hyper-parameter space. As illustrated in Fig 4 and the observations discussed in Section 3, the set of pre-processing hyperparameters affect the performance together, varying across the different datasets. A large area of tissue is detected as background or holes and then dropped during the pre-processing. As illustrated in Fig 4, dropping too many patches decreases performance in Fig. 4(a) and (c) but increases performance in Fig. 4(b) where the dataset is changed. It is hard and no work to measure how these patches and specific types of cells affect the out-of-domain inference performance in histopathological tasks. Thus, we optimize these histopathological tasks by machine learning techniques instead of exploring the impact of the specific type of cells.
As illustrated in Fig 2, the hyper-parameter optimization for histopathological tasks is non-convex optimization which has many large local optima. Therefore, our goal is to search for an optimal hyper-parameter in a short time and with less cost; the optima we search for can just be larger as it can. We are not searching for the global optima. We just search for an optimal value much better than the default hyper-parameter in a very short time and cost. In practice, this involves repeatedly applying different features within the same foundational model using various hyperparameters $H_{n}^{*}$. Therefore, we propose a PSNR mechanism as a threshold to advance the speed of hyper-parameter search.
4.3 Optimization of Fast Hyper-Parameter Search
4.3.1 Good Pre-Processed WSIs show High Structural Similarity.
Figure 5: Relationship between PSNR and Accuracy. The red star stands for the reference object. All the hyper-parameters are compared with the reference object to calculate the PSNR.
For single-domain inference performance, our experiments demonstrate that pre-processing parameters, which lead to similar high accuracy, also generate the pre-processed images with similar structures(as illustrated in Figure 5). When the two objects are close in accuracy, their PSNR is always high. This phenomenon always exists, no matter how we change the reference object.